problem 1: A boy has P 116 annual income from his two savings accounts bearing 3 percent and 5 percent interest. Then he has 25 percent more of the 3 percent savings and 40 percent more of the 5 percent savings, thereby increasing his annual income by P 41. Find his initial investment in each type of savings.
pls. help me solve this problem. tnx

Answer by mananth(16949)

let investment in 3 % be x
investment in 5% be y
...
0.03x+0.05y= 116
3x+5y=11600.............................1
....
25% more of 3% = 1.25x
40% more of 5% = 1.40 y
..
0.03*1.25x+0.05*1.40y=116+41
0.0375x+0.0700y= 157
375x+700y=1570000.......................2
multiply equation 1 by -140 and add it to equation 2
-420x-700y=-1624000
-45x=-54000
x =1200 @ 3%
plug value of x in equation 1
3*1200+5y=11600
5y=11600-3600
5y=8000
y=1600 @5%

C H E C K
0.05*1600 + 0.03*1200=116
aaaaaaah